OpenStudy (anonymous):

Factor k2 + 13k + 12.

OpenStudy (apoorvk):

13k hmm... we need to split that right?

OpenStudy (anonymous):

i think

OpenStudy (apoorvk):

I see a 12 to the right, and a '1' to the left (with k^2), and merlin's beard, 12 +1 do make 13!!! what do you think now? :p

OpenStudy (apoorvk):

*does

OpenStudy (anonymous):

just to tell you the two is a exponent

OpenStudy (apoorvk):

Yeah I understand that.. umm... \[k^2 + 13k + 12=0\] Now, I split '13k' as sum of '12k' and 'k', suits you?

OpenStudy (anonymous):

\[k^2 + 13k + 12 \] \[= k^2 + k + 12k + 12 \] \[= k(k+1)+12(k+1) \] \[= (k+1)(k+12)\]